OpenStudy (anonymous):

How do I set up: over the last 3 years, salesperson A has earned twice as much as salesper B who in turn has averaged 50% more than salesper C. The 3 combined to make $220,000 ver 3 years. How much did salesper B earn on average?

OpenStudy (anonymous):

60000

OpenStudy (anonymous):

A has earned twice as much as B...so A=2B B has averaged 50% more than C...so B=C+0.5C Combined is 220K...So A+B+C=220,000 The last equation can be re-written as 2B+1.5C+C=220K AND since A=2B it can be rewritten again as 3C+1.5C+C=220K now you can get C finally, since B=1.5C you can get B, then to get the average, just divide that by 3 (years) and you should be done.

OpenStudy (anonymous):

60000 for 3 years to B